Accelerator Update for March 30, 2009 - April 1, 2009 |
Monday March 30 The day shift began with Operations monitoring Tevatron (TeV) store 6943 (74.12E30) and the Recycler stash (266.2E10), with the Antiproton Source stacking (14.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. At 9:27 AM, Operations reported a LCW leak from a power supply (V901A) at MI-62. A power supply technician investigated. At 10:11 AM, a Linac RF station (LRF2) tripped off for the first of two times in two minutes. It reset. Operations terminated store 6943 at 11:03 AM. At 11:57 AM, LRF4 tripped off. It reset. At 12:05 PM, Operations reported that the 400 MeV area beam permit had tripped off for the first of three times in five minutes. Operations put Pbar into standby at 12:22 PM and turned off the MI beam and ramp. This was done so that Electrical Engineering Support techs could search for a possible ground fault on the QF bus between MI-40 and MI-60. They did not find a fault. Operations established store 6944 at 12:33 PM with an initial luminosity of 303.48E30. MI resumed sending beam to NuMI and Pbar at 12:41 PM. Operations reported at 3:57 PM that the Recycler had stashed antiprotons seven times during the day shift. A MI Physicist began a 10-minute study at 4:06 PM. At 4:42 PM, the 400 MeV area beam permit tripped off for the first of two times in five minutes. It reset. At 7:25 PM, the NuMI beam permit tripped off. It reset. At 7:56 PM, a Linac Klystron RF station (KRF6) tripped off for the first of two times in 16 minutes. It reset. At 7:59 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 and 14 to also trip. This happed twice in 15 minutes. Everything reset. At 8:39 PM, the MI South Anode power supply tripped off and brought down a number of MI RF stations. Everything reset. At 8:50 PM, LRF3 tripped off. It reset. Operations reported at 11:58 PM that the Recycler had stashed eight times during the evening shift. Tuesday March 31 The midnight shift began with Operations monitoring store 6944 (90.39E30) and the stash (289.1E10), with Pbar stacking (25.7E10), and with NuMI and MiniBooNE taking beam. At 12:33 AM, LRF3 tripped off. It reset. At 12:38 AM, a Booster vacuum pump (IPRF11) tripped off and caused BRF9 and 10 to also trip. Everything reset. At 1:15 AM, BRF11 tripped off. It reset. At 1:46 AM, LRF4 tripped off. It reset. At 1:49 AM, the 400 MeV area tripped off. It reset. At 2:04 AM, MIRF4 tripped off and would not reset. Operators investigated. At 2:18 AM, the NuMI beam permit tripped off when a power supply (E:V108) tripped off. Everything reset. At 2:36 AM, the low energy vacuum pump for LRF3 tripped off. It reset. At 4:20 AM, LRF3 tripped off. It reset. Operations terminated store 6944 at 6:35 AM. Cryo system technicians began work at TeV sectors A3 and E2 at 6:43 AM. Operations prepared D0 for an access at 6:46 AM. The Neutron Therapy Facility (NTF) reported at 7:18 AM that they had begun treatment on their first of three patients for the day. At 7:37 AM, the NuMI beam permit tripped off on a target position indication, but this trip was preceded by an I- Source spark, which might have been the cause. Operations reported at 8:07 AM that the Recycler had stashed seven times during the midnight shift. D0 came out of access at 8:19 AM. Operations reported at 8:21 AM that the Recycler set a record with a stash size of 477.2E10. At 9:03 AM, the 400 MeV area tripped off. It reset. Cryo system technicians reported at 10:24 AM that they had completed their wet engine work at TeV sectors A3 and E2. At 10:34 AM, a TeV power supply (D0Q5) tripped off on a lead failure. The Operations TeV Specialist adjusted the lead limits. Operations put the TeV through a dry squeeze at 11:01 AM. NTF reported that they had completed treatment on their last patient for the day at 11:10 AM. Operations put the TeV through a wet squeeze at 11:45 AM. Operations put the TeV through another wet squeeze at 12:24 PM. Operations put the TeV into shot setup at 12:46 PM. Operations reported at 1:12 PM that the Recycler stash had been partially mined for the next store. Operations established store 6949 at 2:10 PM with an initial luminosity of 331.98E30. At 3:45 PM, LRF5 tripped off. It reset. Operations reported at 3:57 PM that the Recycler had stashed three times during the day shift. TeV Physicists reported at 4:27 PM that they had moved the crystal collimator to a position approximately one centimeter away from the beam. At 10:19 PM, the NuMI beam permit tripped off due to a MI RF indication, but operators checked and there was nothing wrong with the MI RF system. The beam permit reset. At 11:23 PM, LRF4 tripped off. It reset. Operations reported at midnight that the Recycler had stashed nine times during the evening shift. Wednesday April 1 The midnight shift began with Operations monitoring store 6949 and the stash (412.4E10), with Pbar stacking (26.0E10), and with NuMI and MiniBooNE taking beam. Operations terminated store 6949 at 12:08 AM. At 12:32 AM, a TeV kicker (C:AAK1) tripped off on a “prefire/fail to fire” indication. It reset. Operations established store 6950 at 1:43 AM with an initial luminosity of 350.23E30. NTF reported at 7:12 AM that they were beginning treatment on their first of three patients for the day. At 7:51 AM, four Linac quadrupole power supplies (L:QPS104, 113, 114, and 128) tripped off. The Linac Group Leader and operators responded. They got all of the power supplies back on except for 128. The Linac Group Leader investigated further. Operations reported at 8:01 AM that the Recycler had stashed six times during the midnight shift. The Plans for Wednesday The plans for today are to stack, stash, and store. Store 6950 will be terminated around 5 PM and D0 and CDF will be allowed to access. |
|
More Information Comments and Suggestions |
| last modified 04/01/2009 email Fermilab |
| Security, Privacy, Legal | |